Consolidated Industries is a diversified manufacturer with business units organized as divisions, including the Reigis Steelper unit 1. Contribution margin 2a. Return on investment 2b. Residual income
Answer #1

Answer:

Contribution margin = 56000-29000-(4500/3) = 25500

a) Contribution margin per unit = 25500000/1700000 = 15 per unit

b) ROI = Net income/Operating assets = 16500000/114250000= 14.44%

c) Residual income = 16500000-(114250000*10%) = 5075000